# The Valuable Long-period Cluster Cepheid KQ Scorpii and other Calibration Candidates

### v1

@@ 2 Analysis
Fig. 1 highlights stars with proper motions ($\mu_{\alpha}$, $\mu_{\delta}$) within $0.2$ mas yr-1 of KQ Sco (Gaia Collaboration et al., 2023). An overdensity is readily discernible which represents UBC 1558, and KQ Sco is within the confines of the cluster (Fig. 1). UBC 1558 is not densely populated, and may not remain bound as the average cluster dissolution timescale is surpassed (10 Myr, Bonatto & Bica, 2011). The Gaia DR3 astrometric solutions for the cluster and KQ Sco are summarized in Table 1. The parallaxes agree to within the uncertainties. A Wesenheit function (Leavitt Law) was employed to independently assess the distance to KQ Sco. The compilation of Ngeow (2012) presents a Cepheid Wesenheit distance of $\simeq 2.2$ kpc, and the relation of Majaess et al. (2013a) corroborates the result. Those distances are nearer, thereby hinting at a correction to the DR3 results (e.g., Lindegren et al., 2021; Owens et al., 2022). Applying the Lindegren et al. (2021) prescription bridges the gap, and suggests a mean systematic correction of $\Delta\pi\simeq 0.03$ mas for these DR3 results. KQ Sco shifts from 0.43 to 0.47 mas, which is a $\simeq 10$% distance offset (pc).
@@
Table 2: Understudied Cluster Cepheid candidates warranting individual follow-up.
@@
CV Mon F 5.4 ASAS J151832.37-580128.7 F 9.2
DP Vel F 5.5 ASAS J183904-1049.3 1O 3.1
EX Vel F 13.2 ASAS J192152.00+150346.9 F 8.8
FF Car F 16.3 ASAS J194806.54+260526.1 1O 6.6
FM Car F 7.6 ASAS J201151.18+342447.2 F 9.8
FZ Car 1O 3.6 ASAS J211659.94+514556.7 F 5.9
GI Cyg F 5.8 ASAS J213533.70+533049.3 1O 3.2
@@ 3 Conclusions
KQ Sco exhibits parameters which are consistent with UBC 1558, and the Cepheid is a probable member (Fig. 1, Table 1). The distance to UBC 1558 was revised nearer by $\simeq-0.8$ kpc. An unweighted mean and standard deviation associated with all distance methods yield $d=2.15\pm 0.15$ kpc (excluding uncorrected DR3 parallaxes).
Going forward, numerous understudied potential Cepheids may be cluster members (Table 2). The targets were identified by cross-referencing the Pietrukowicz et al. (2021) Cepheid catalog with DR3. A subsample were discussed previously (e.g., Zhou & Chen, 2021; Lin et al., 2022), and all candidates shall benefit from individual follow-up, which presents pertinent insights as conveyed here. Owing to the importance attributed to cluster Cepheids in diverse endeavors (e.g., ascertaining whether the Planck $\Lambda$CDM $H_{0}$ requires adjustment): it is desirable that a site be constructed to provide suggested guidance for the broader community whereupon (I), an updated list of cluster Cepheids unanimously agreed upon by a panel of international researchers is presented, and (II), reputed cluster Cepheids lacking consensus are highlighted which may require additional funding to secure data and undertake a comprehensive analysis. For example, QZ Nor (Eggen, 1983; Majaess et al., 2013b) is featured in the Gold sample of Reyes & Anderson (2023) (see also Breuval et al., 2020), but absent from Medina et al. (2021) and Hao et al. (2022). Conversely, GQ Vul is classified within the Class A sample of Hao et al. (2022), and is absent from Reyes & Anderson (2023), and a 0.47 membership probability was assigned by Medina et al. (2021, their Table 1). Turner et al. (2008) suggest CG Cas is a member of Berkeley 58, whereas Wang et al. (2024) indicate it is a member of NGC 7790. Therein lies the motivation for researchers to create and frequently update a cluster Cepheid site.
Also looking forward on a separate front, an investigation is desirable to examine whether bound or dissolving clusters encompassing UBC 1558 (Fig. 1) are associated (e.g., Ruprecht 121). For example, the denser open cluster NGC 6216 may be in relative vicinity to UBC 1558, yet the former may be potentially older and therefore unrelated.

### v2

@@ 2 Analysis
Fig. 1 highlights stars with proper motions ($\mu_{\alpha}$, $\mu_{\delta}$) within $0.2$ mas yr-1 of KQ Sco (Gaia Collaboration et al., 2023). An overdensity is readily discernible which represents UBC 1558, and KQ Sco is within the confines of the cluster (Fig. 1). UBC 1558 is not densely populated, and may not remain bound as the average cluster dissolution timescale is surpassed (10 Myr, Bonatto & Bica, 2011). The Gaia DR3 astrometric solutions for the cluster and KQ Sco are summarized in Table 1. The parallaxes agree to within the uncertainties. A Wesenheit function (Leavitt Law) was employed to independently assess the distance to KQ Sco. The compilation of Ngeow (2012) presents a Cepheid Wesenheit distance of $\simeq 2.2$ kpc, and the relation of Majaess et al. (2013a) corroborates the result. Those distances are nearer, thereby hinting at a correction to the DR3 results (e.g., Lindegren et al., 2021; Owens et al., 2022). Applying the Lindegren et al. (2021) prescription bridges the gap, and suggests a mean systematic correction of $\Delta\pi\simeq 0.03$ mas. For example, KQ Sco shifts from 0.43 to 0.47 mas, which is a $\simeq 10$% distance offset (pc).
@@
Table 2: Understudied cluster Cepheid candidates warranting follow-up.
@@
CV Mon F 5.4 ASAS J183904-1049.3 1O 3.1
DP Vel F 5.5 ASASSN-V J040516.13+555512.9 1O 1.8
EX Vel F 13.2 ASASSN-V J151832.37-580128.7 F 9.2
FF Car F 16.3 ASASSN-V J194806.54+260526.1 1O 6.6
FM Car F 7.6 ASASSN-V J201151.18+342447.2 F 9.8
FZ Car 1O 3.6 ASASSN-V J211659.94+514556.7 F 5.9
GI Cyg F 5.8 ASASSN-V J213533.70+533049.3 1O 3.2
@@ 3 Conclusions
KQ Sco exhibits parameters which are consistent with UBC 1558, and the Cepheid is a probable member (Fig. 1, Table 1). The distance to UBC 1558 was revised nearer by $\simeq-0.8$ kpc. An unweighted mean and standard deviation associated with all distance methods yield $d=2.15\pm 0.15$ kpc (excluding uncorrected DR3 parallaxes). A separate investigation is desirable to examine whether bound or dissolving clusters encompassing UBC 1558 (Fig. 1) are associated (e.g., Ruprecht 121). For example, the denser open cluster NGC 6216 may be in relative vicinity to UBC 1558, yet the former may be potentially older and therefore unrelated.
Going forward, numerous understudied potential Cepheids may be cluster members (Table 2). The targets were identified by cross-referencing the Pietrukowicz et al. (2021) Cepheid catalog with DR3. A subsample were discussed previously (e.g., Zhou & Chen, 2021; Lin et al., 2022), and all candidates may benefit from individual follow-up, which presents pertinent insights as conveyed here.
More broadly, owing to the importance attributed to cluster Cepheids in diverse endeavors (e.g., ascertaining whether the Planck $\Lambda$CDM $H_{0}$ requires adjustment): a database site could be constructed to provide suggested guidance for the broader community whereupon (I), an updated list of cluster Cepheids unanimously agreed upon by a panel of international researchers is presented, and (II), reputed cluster Cepheids lacking consensus are highlighted which may require additional funding to secure data and undertake a comprehensive analysis. For example, QZ Nor (Eggen, 1983; Majaess et al., 2013b) is featured in the Gold sample of Reyes & Anderson (2023) (see also Breuval et al., 2020), but absent from Medina et al. (2021) and Hao et al. (2022). Conversely, GQ Vul is classified within the Class A sample of Hao et al. (2022), and is absent from Reyes & Anderson (2023), and a 0.47 membership probability was assigned by Medina et al. (2021, their Table 1). Turner et al. (2008) suggest CG Cas is a member of Berkeley 58, whereas Wang et al. (2024) indicate it is a member of NGC 7790. Therein lies the motivation for researchers to maintain a cluster Cepheid site.